If the demand function faced by a firm is:

Q = 90 – 2P

TC = 2 + 57Q – 8Q2 + Q3


Determine the level of output at which the firm maximizes the profit.

Determine the best level of output for the above question by the MR and MC approach.



1
Expert's answer
2021-05-06T14:30:56-0400

"TP_L = 10L + 5L^2 + L^3"


"MP_L=\frac {dTP_L} {dL} =10+10L+3L^2"


"AP_L=\frac {TP_L} {L} =10+5L+L^2"



When L=5;


"TP_L=10(5) + 5(5)^2 + (5)^3"


"TP_L=50+125+125"


"TP_L=300"



"MP_L=10+10(5)+3(5)^2"


"MP_L=10+50+75"


"MP_L=135"



"AP_L=10+5(5)+(5)^2"


"AP_L=60"